OpenStudy (anonymous):

The numbers of beads on 500 handcrafted bead necklaces follow a normal distribution whose mean is 24 beads and standard deviation is 4 beads. Which sentence most closely summarizes the data?

OpenStudy (anonymous):

About 80 necklaces have more than 28 beads. Z score is 28-24/4=1 (1-.8413)=0.1587 500*0.1587=79.35~80 hope this helps :)